Opened 5 years ago

Closed 3 years ago

#553 closed flight processing (fixed)

RG13/06, flight day 298/2014, Danum Valley

Reported by: dap Owned by:
Priority: alpha 5 Milestone:
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by benj)

Data location: /users/rsg/arsf/arsf_data/2014/flight_data/malaysia/RG13_06-2014_298_Danum_Valley

Data arrived from network transfer on 17/12/2014.

Scientific objective: Investigating the impact of humans on tropical forests.

Priority: alpha 5

PI: David Coomes

No owl data present

Corrupt webcam images were removed:
141024_232542_000.jpg

Sensors:

  • Fenix (requested)
  • Leica FW LIDAR (?)
  • Leica LIDAR (requested)
  • RCD (requested)

Duplicate lines

Process only one for Fenix and RCD - check for which has least cloud cover, failing that choose closest to local noon

  • 1188-1200 (day 295, 1188-1189 repeated 3 times, also day 294)
  • 1-9

Change History (41)

comment:1 Changed 5 years ago by dap

  • Priority changed from immediate to alpha 5

comment:2 Changed 4 years ago by benj

  • Description modified (diff)

comment:3 Changed 4 years ago by tec

Nav
Data in ~/arsf_data/2014/flight_data/malaysia/RG13_06-Basestation_data/RINEX DATA SABAH/Rinex/temporary_gps_dataz/298, move out into project folder once processed.

comment:4 Changed 4 years ago by benj

  • Description modified (diff)

Re-flown on day 100 (#572). Not processing this unless additional funding made available.
Ignore me, browser fail...

Last edited 4 years ago by benj (previous) (diff)

comment:5 Changed 4 years ago by benj

  • Description modified (diff)

comment:6 Changed 4 years ago by lah

Navigation Processing

Latitude4 43 30.55967
Longitude117 36 19.18848
El Height480.125 m
Last edited 4 years ago by dac (previous) (diff)

comment:7 Changed 4 years ago by dac

Navigation Processing

  • Reprocessed basestation data using correct LEAS10 profile (had used incorrect LEAR25 previously)
  • Still couldn't get good result in IPAS Pro but good result obtained in IPAS TC with default settings.
  • Processed files 2 and 3 separately and combined to create single sol file.
Last edited 4 years ago by dac (previous) (diff)

comment:8 Changed 4 years ago by lah

Navigation Processing
Created new combined sol file using all 3 files to stop alspp errors (20141025_lah.sol)

comment:9 Changed 4 years ago by lah

Navigation Processing
Have redone the camera sol file as well given multiple requests. Previous version and separate files moved to subdirectory.

comment:10 Changed 4 years ago by lah

Lidar Processing
Found roll and pitch for all flightlines: r= -0.0062, p= -0.0006.
Some lines are extremely short, so may not be worth keeping.

comment:11 Changed 4 years ago by tec

RCD Processing
Starting

comment:12 Changed 4 years ago by lah

Fenix processing started
Lines 14-19 vnir binning = 4 2, all other lines 2 2 as normal.

comment:13 Changed 4 years ago by tec

RCD Processing
Removing images 137 261 349 350 381 382 387 388 389 395 396 415 416 417 418 419 420 423 424 425 426 430 431 432 434 435 436 437 438 439 446 447 448 449 450 451 452 453 454 457 460 461 462 463 464 465 466 467 468 469 470 475 476 478 481 482 483 484 485 489 499 500 501 502 504 505 506 508 509 510 511 512 513 515 518 519 520

comment:14 Changed 4 years ago by lah

Fenix Processing
Not processing lines 11 and 12 (1189 & 1188) as there are less cloudy duplicates in day 295. 1190 doesn't appear to have been saved or has been overwritten (crossed out in logsheet as another line 11). 1191 - 1200 (lines 10-1) are less cloudy for 298 than 295, so will be processed. 1-9 have not been duplicated in 298.

Last edited 4 years ago by lah (previous) (diff)

comment:15 Changed 4 years ago by lah

Fenix Processing
Found SCTs:

Flightline FENIX
1 1.00
2 0.94
3 0.98
4 0.98
5 1.00
6 0.00
7 1.01
8 0.98
9 1.01
10 0.92
13 0.96
14 0.00
15 0.90
16 0.98
17 0.96
18 0.98
19 0.99
20 1.06
21 0.90
22 0.00
23 0.99
24 0.96
25 0.98
26 0.94
27 1.01
28 0.93
29 0.99
30 0.99
31 0.00
32 0.97
33 0.97
34 0.96
35 1.00
36 0.93
37 1.00
38 0.96
39 0.97
40 0.00
41 1.03
42 0.90
43 1.03
44 0.96
45 0.99

comment:16 Changed 4 years ago by dac

LiDAR Processing

  • Fixed problem where LAS files had a large offset set for y in header. Set all values to 0 using laspy (same as for other files). Checked only offset and min/max y were changed using lasdiff.
  • Recreated ASCII files.
  • Generated DEMs, now sensible size and lidar lines up OK.

comment:17 Changed 4 years ago by lah

Fenix ready for DC

Line 006a doesn't have any fenix data (and no file number) File 20 does exist, but is not mentioned on the logsheet. Have noted this in the readme.

comment:18 Changed 4 years ago by lah

Lidar ready for DC

001f could not be processed by aslpp, so is not included. Have mentioned in readme.

comment:19 Changed 4 years ago by dap

Beginning Fenix Delivery Check

comment:20 Changed 4 years ago by tec

LiDAR DC
Started

comment:21 Changed 4 years ago by dap

Fenix Delivery Check

  • Ran rm *.aux.xml from flightlines/mapped to remove some xml files that had been left in my tuiview.
  • Ran mv fenix_mosiac.jpg fenix_mosaic.jpg in screenshots directory to rename fenix mosaic
  • Looking at screenshots 19, 20 and 29, these flightlines look like they're oddly shaped - do you know why this is? Is it worth commenting on in the read me?
  • Duplicate lines have been processed (1164 - 1184, files 21 - 42) - the duplicates from Julian day 293 were processed and have less cloud coverage. I've spoken to benj and we don't need to take these out, so duplicates aren't a problem.

comment:22 Changed 4 years ago by tec

LiDAR DC
Demcompare says 18.4, assuming this is find the delivery is ok.

comment:23 Changed 4 years ago by lah

Fenix

Have modified readme to explain oddly shaped lines.

comment:24 Changed 4 years ago by dap

Fenix Delivery Check

New data quality remarks look fine, but over-/underflows table has been pushed to the next page, separate from the section header, which looks a bit odd, so I've put it on the next page with the table.

Fenix dataset ready for delivery.

Last edited 4 years ago by dap (previous) (diff)

comment:25 Changed 4 years ago by lah

Fenix & Lidar Rsynced to ftp (arsf1)

comment:26 Changed 4 years ago by tec

RCD Processing
Tagging tiffs. Removed the first ~7 from the csv file as they were causing errors to do with the sol file, they were not part of any line so this is fine.

comment:27 Changed 4 years ago by tec

RCD Processing
Ready for DC

comment:28 Changed 4 years ago by dac

RCD DC

Starting delivery check.

comment:29 Changed 4 years ago by dac

RCD DC

Looks fine - ready to deliver.

Events file only had read permission as airborne - have fixed

comment:30 Changed 4 years ago by lah

Last edited 4 years ago by lah (previous) (diff)

comment:31 Changed 4 years ago by lah

RCD uploaded to ftp and email sent 12/06/15

comment:32 Changed 4 years ago by stgo

Starting Fenix Reprocessing

comment:33 Changed 4 years ago by stgo

Fenix reprocessing

Ready for delivery check.

This delivery needs to be thoroughly checked for:

  • correct lines included
  • correct linenames from logsheet
  • data quality comments
  • spectra shifts

comment:34 Changed 4 years ago by gej

Fenix DC

Files 20,24,29,39 and 43 are weirdly shaped. Logsheet seems to suggest this is due to an F2 error. Needs to be in readme

Screenshots also need renaming.
Logsheet needs renaming according to proj_tidy
Data quality report needs changing to 2015 version.
SCT's look okay.

Last edited 4 years ago by gej (previous) (diff)

comment:35 Changed 4 years ago by lah

Fenix DC - spectra
Checked spectra of all lines and no major problems. Most lines have a spike at the end of the VNIR and another at the start of the SWIR, within the mask region. The spectral shift looks small and is within 5 nm compared to previous data (2014 169).

comment:36 Changed 4 years ago by stgo

Fenix DC

I've updated the readme with a comment about the weirdly shaped lines and a warning about potential effects on data.

Left screenshots for gareth to test make_arsf_delivery changes on, please back up the screenshots!

Data quality report updated.

comment:37 Changed 4 years ago by gej

Fenix DC

Readme comments are better,
Screenshots have been renamed, proj_tidy.sh thinks logsheet is incorrectly named, but I think it's okay as it is due to it being in the SAFE area.

Data quality report is correct

Zipping mapped files now.

comment:38 Changed 4 years ago by gej

Fenix Delivered to Dr. Coomes 16th november 2015 via FTP on user arsf1

comment:39 Changed 3 years ago by dac

Archiving

Checked and ready for upload to NEODC.

comment:40 Changed 3 years ago by dac

Archiving

Uploaded to NEODC.

comment:41 Changed 3 years ago by asm

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.